During a nine hour snow storm, it snows at a rate of one inch per hour for the first two hours, at a rate of 2 inches per hour for the next 6 hours, and at a rate of 0.5 inches for the remaining hour. Write a piecewise-defined function that gives the depth of the snow storm. How many inches of snow accumulated from the snow storm?
Answer by Theo(13342)

let t = total inches of snow and let h = number of hours that it snowed.

then:

t = 1 * h for h <= 2 + 2 * h for 2 < h <= 8 + .5 * h for 8 < h <= 9

total accumulation of snow would be 1 * 2 + 2 * 6 + .5 * 1 = 14.5 inches of snow.